You mix 100.0 mL of 0.20 M HBr and 50.0 mL of 0.40 M NaClO. What is the pH of the resulting solution? Ka(HClO) = 3.5 x 10^-8. Provide your answer with solution.


1
Expert's answer
2021-04-16T04:59:40-0400

Balanced equation for the reaction;

2HBr + NaClO "\to" Br2 + NaCl + H2O


Moles of HBr


If 0.2 moles were in 1000mL,

How many will be in 100mL


"=\dfrac{0.2 mol x 100mL}{1000mL} =" 0.02 mol


Moles of NaClO that is required to completely react with HBr;


Mole ratio of HBr : NaClO is 2:1


Moles of NaClO "=\dfrac{1}{2}x0.02 = 0.01mol"


But moles of NaClO present were;


0.4 mol were in 1000mL

How many will be in 50mL


"=\dfrac{0.4mol x 50mL}{1000mL} = 0.02mol"


Thus moles of NaClO are in excess by 0.02 mol - 0.01 mol = 0.01 mol


Therefore we are actually determining the pH of 0.01mol NaClO in 150mL of solution (since all the other products are neutral)


Concentration of the remaining NaClO;


0.01 mol were in 150mL of solution (100+50)

How many will be in 1000mL


"=\dfrac{0.01mol x 1000mL}{150mL} = 0.0667 M"


We then determine pH of 0.0667M NaClO


The salt will ionize as; (since Na+ is neutral)


ClO- + H2O "\iff" HClO + OH- (Kb = 3.45 x 10-7)


I 0.0667 0 0

C -X +X +X

E 0.0667-X X X


"Kb = \dfrac{[HClO][OH]}{[ClO]}"


"0.0667x3.45x10^-7 = \dfrac{X^2}{0.0667} x 0.0667"



"0.0667x3.45x10^-7 = X^2"


finding squareroot of both sides

"\surd0.0667x3.45x10^-7 = X"


X = 4.7958x10-4 = [OH]


pOH = -Log [OH]

= -Log 4.7958x10-4

= 3.319


Since pOH + pH = 14

pH = 14-pOH

=14-3.319

= 10.68

Answer: pH = 10.68